# Understanding Action Effects through Instrumental Empowerment in Multi-Agent Reinforcement Learning

### v1

@@ Abstract
To reliably deploy Multi-Agent Reinforcement Learning (MARL) systems, it is crucial to understand individual agent behaviors within a team. While prior work typically evaluates overall team performance based on explicit reward signals or learned value functions, it is unclear how to infer agent contributions in the absence of any value feedback. In this work, we investigate whether meaningful insights into agent behaviors can be extracted that are consistent with the underlying value functions, solely by analyzing the policy distribution. Inspired by the phenomenon that intelligent agents tend to pursue convergent instrumental values, which generally increase the likelihood of task success, we introduce Intended Cooperation Values (ICVs), a method based on information-theoretic Shapley values for quantifying each agent’s causal influence on their co-players’ instrumental empowerment. Specifically, ICVs measure an agent’s action effect on its teammates’ policies by assessing their decision uncertainty and preference alignment. The analysis across cooperative and competitive MARL environments reveals the extent to which agents adopt similar or diverse strategies. By comparing action effects between policies and value functions, our method identifies which agent behaviors are beneficial to team success, either by fostering deterministic decisions or by preserving flexibility for future action choices. Our proposed method offers novel insights into cooperation dynamics and enhances explainability in MARL systems.
@@ 1 Introduction
In MARL, the reward function is fundamental in shaping agent behavior, with Schmidhuber’s artificial curiosity [33] inspiring the design of intrinsic rewards to foster cooperation. As the majority of MARL algorithms (see, e.g., [30]) rely on learned value functions to estimate expected rewards in a given state, the SV naturally emerges as a credit assignment method in fully cooperative MARL scenarios. The credit assignment problem in MARL concerns identifying each agent’s contribution to the team’s overall performance. Conversely, in the absence of both value functions and reward signals, a critical question arises of how, and more importantly which, credit should be allocated, irrespective of the game type (cooperative or competitive). In this setting of missing explicit value feedback, we investigate whether SVs can be adapted to extract meaningful information about agent behavior on performance, i.e., information that remains consistent with the underlying value function, while assuming access only to the policy function.
To identify a meaningful candidate for credit, we take inspiration from the famous statement by von Foerster [41]: "I shall act always so as to increase the total number of choices", thereby advocating for individual empowerment. In a related vein, Omohundro [28] derived the fundamental tendency of optimal agents to seek power. Both philosophies align with the instrumental convergence thesis by Bostrom [5], which posits that several instrumental values are universally beneficial to increase task success and thus likely to be pursued by many intelligent agents. It has been theoretically [4, 39] and empirically [10, 20] shown that access to and control over a wide range of future states (or options) serves as a proxy for preparedness. This makes it a convergently instrumental objective, as formally defined by the empowerment function [16] or by the policy entropy [14]. Given the evidence that state space exploration is generally beneficial during training, we question whether the same holds true during inference in multi-agent settings. Concretely, is providing agents with more options truly beneficial for task success?
@@
Figure 1: Left: a toy environment where both players must reach the target (star) to earn a shared reward, while gray cells block the path. Blue and red arrows indicate potential moves (choices). Right: probabilities (y-axis) over actions (x-axis) of P2 before (top) and after (bottom) P1 steps on the key cell and opens the green lock. By opening the green key, P2 becomes more certain about choosing the leftward action, whereas unlocking the gray locks would provide additional movement options that are clearly not beneficial for P2.
@@
Drawing on the classic exploration-exploitation dilemma in Reinforcement Learning (RL) [37], we argue that exploiting certain situations with high certainty may be more instrumentally valuable than preserving many choices. We illustrate this concept in Fig. 1. The maze on the left shows a toy environment where two agents must reach a target (star) to receive a shared reward of one. The histograms on the right depict the probability distributions over movements actions for P2 before (top) and after (bottom) P1 acted. Consider the case where stepping on the key opens the green lock only. In the current state, the lock cell to the left of P2 is closed, which makes P2 more likely to move downward, while still assigning some probability to the leftward action in anticipation that P1 might open the lock. Once P1 steps on the key cell, this action causally affects P2’s action space by unlocking the left path. As a result, P2 not only gets more movement options, but also becomes more certain in exploiting the open path by assigning more weight toward the left action. This highlights how P2’s (increased) success is directly tied to P1’s altruistic behavior, which would correspond to an increased value under an optimal value function. In contrast, if the key were to unlock the gray locks, it would provide additional movement options that are clearly not beneficial for achieving its goal. Even if the game involves simultaneous decision-making and P2 chooses to go down, P1 would merit credit for his intention to open the path and contribute to P2’s success. In addition to analyzing effects on decision uncertainty, we examine the impact on preference alignment between agents. To do so, we perform a state exchange intervention by placing P1 in P2’s position and v.v. to assess whether agents exhibit coordinated consensus through mutual anticipation of similar behavior or whether they favor divergent strategies. Credit is attributed to actions that enhance alignment and promote consensus.
@@
Motivated by this, we introduce a causal action attribution method called Intended Cooperation Value (ICV), which quantifies how cooperatively an agent behaves by measuring the causal effect of its actions on the instrumental empowerment of its teammates, using only access to their policies. By adapting causal, information-theoretic Shapley Values, ICV uses the entropy of the policy distributions to assign credit to an agent based on its intended contribution to reducing decision uncertainty and fostering preference alignment among co-players. Empirical evaluations across cooperative, competitive, and mixed-motive settings demonstrate that ICV reliably attributes credit to behaviors that increase the likelihood of task success and aligns with standard value-based SVs. Our approach provides a deeper understanding of agent interactions, particularly in scenarios where success is not explicitly measurable.
2 Background
@@
Let $(\Omega,\mathcal{F},\Pr)$ be a probability space. Let $X:\Omega\rightarrow\mathcal{X}$ be a discrete Random Variable (RV) taking value $x$ in a set $\mathcal{X}$, where $p(x)=\Pr(X=x)$ is the Probability Mass Function (PMF) and $\mathcal{P}(\mathcal{X})$ is the set of probability distributions over $\mathcal{X}$. We restrict our attention to finite sets $\mathcal{X}$. The entropy of $X$ is given by $H(X)=-\sum_{x\in\mathcal{X}}p(x)\log p(x)$. For two PMFs $p$ and $q$, the Kullback–Leibler divergence is $D_{KL}(p||q)=\sum_{x\in\mathcal{X}}p(x)\log\frac{p(x)}{q(x)}$.
2.1 Reinforcement Learning
@@
The interactions among $n$ concurrently acting agents are typically modeled using either a Markov Game (MG) [22] or an Extensive-Form Game (EFG) [29]. While in EFGs actions are selected and executed sequentially, in MGs actions are selected simultaneously but their execution may be either parallel or sequential, depending on whether the environment supports parallelization. If not, actions are executed based on a predefined or randomized agent ordering.
@@
Assumption 1 implies that we consider domains with full state observability, though our approach can trivially be extended to partial observability and sequential decisions.
@@
An MG is defined by the tuple
@@
where $\mathcal{N}=\{1,\dots,n\}$, with $n\in\mathbb{N}$, is the set of agents, $\mathcal{S}$ is the set of admissible states, $\mathcal{A}^{i}$ is the set of actions available to agent $i\in\mathcal{N}$, with $\boldsymbol{\mathcal{A}}=\prod_{i=1}^{n}\mathcal{A}^{i}$ being the joint action set, $P:\mathcal{S}\times\boldsymbol{\mathcal{A}}\to\mathcal{P}(\mathcal{S})$ is the state transition probability kernel, $\rho\in\mathcal{P}(\mathcal{S})$ is the initial state distribution, $R^{i}:\mathcal{S}\times\boldsymbol{\mathcal{A}}\times\mathcal{S}\to\mathbb{R}$ is the reward function of agent $i$, and $\gamma\in[0,1)$ is the discount factor controlling the effect of future rewards. At each time step $t\in\mathbb{N}$, let the RV $S_{t}$ describe the state of the Markov game, and the RV $A_{t}^{i}$ denote the action taken by agent $i$. Starting from an initial state $s_{0}\sim\rho$, all agents observe the current state $s_{t}\in\mathcal{S}$ and simultaneously select an action $a_{t}^{i}\in\mathcal{A}^{i}$ according to their individual policies $\pi^{i}(a\mid s_{t})$, forming a joint action $\boldsymbol{a}_{t}=(a_{t}^{1},\dots,a_{t}^{n})\in\boldsymbol{\mathcal{A}}$ drawn from the joint policy $\boldsymbol{\pi}(\boldsymbol{a}\mid s_{t})=\prod_{i=1}^{n}\pi^{i}(a\mid s_{t})$. The MG then transitions to a new state $s_{t+1}$ according to $P(s_{t+1}\mid s_{t},\boldsymbol{a}_{t})$, and each agent $i$ receives a reward $R^{i}(s_{t},\boldsymbol{a}_{t},s_{t+1})$. The performance of each agent $i$ can be measured by means of a value function defined as:
$V^{i}(s)\coloneqq\mathbb{E}_{\boldsymbol{\pi},P}\left[\sum_{t=0}^{\infty}\gamma^{t}R^{i}(S_{t},\boldsymbol{A}_{t},S_{t+1})\mid S_{0}=s\right].$ (1)
We restrict our attention to environments satisfying the MG dynamics as per Def. 1. Henceforth we, use the terms environment, game, and MG interchangeably.
@@
Assumption 2.
We consider pre-trained Actor-Critic [17] models under the MG framework using Centralized Training with Decentralized Execution (CTDE) [24], the state-of-the-art paradigm for MARL. Further, we consider stochastic policies, which are generally more robust and tend to perform better in real-world-like environments compared to deterministic policies [46].
From Definition 2, we emphasize the flexibility of MGs to represent all game types [9]. Assumption 2 implies that agents act with individual policies (actors), and in fully cooperative settings, they share a global value function $V(s)$, also called centralized critic [2]. We use $\boldsymbol{-i}$ to denote the set of all agent indices excluding agent $i$.
@@ 2.2 Shapley Value
We build on prior work applying SVs for feature attribution in single-agent RL and Machine Learning (ML) models. Therefore, let $Z$ represent the state in RL or the input in ML.
@@
Let $\mathcal{M}=\{1,\dots,m\}$ be the set of features and $\mathcal{Z}=\prod_{i=1}^{m}\mathcal{Z}_{i}$ the input space, so that any input $z$ can be represented as an ordered set $z=\{z_{i}\mid z_{i}\in\mathcal{Z}_{i}\}_{i=1}^{m}$, with $p(z)$ being the data distribution. Then, any subset $\mathcal{K}\subseteq\mathcal{M}$ yields a partial observation $z_{\mathcal{K}}=\{z_{i}\mid i\in\mathcal{K}\}$, where $\bar{\mathcal{K}}=\mathcal{M}\setminus\mathcal{K}$ is its complement.
@@
The Shapley Value [34] is the unique method that satisfies the desirable fairness properties of efficiency, symmetry, additivity, and null player (see Appendix B). It allocates credit by computing the average marginal contribution $\Delta v(\cdot)$ of player (or feature) $i$ to the outcome of a cooperative game (or prediction) over all coalitions $\mathcal{K}\subseteq\mathcal{M}$ using a characteristic function $v:2^{\mathcal{M}}\rightarrow\mathbb{R}$, as follows:
@@
3 Related Work
There has been considerable work on understanding and explaining RL, primarily aimed at answering why specific actions were taken. Existing approaches include replacing the policy network with transparent models, such as decision trees [38], or incorporating auxiliary reasoning mechanisms like counterfactual analysis [45], structural causal models [26], or interpretable reward components [23]. The SV has been used as a feature attribution method in Machine Learning (ML) [12, 47], in single-agent RL [3, 35], and in MARL to assess agent contributions based on rewards [27]. Beyond explanatory purposes, SVs have also been adapted as training mechanisms in MARL, mainly to address the credit assignment problem [7, 43, 42]. In parallel, various information-theoretic intrinsic rewards have been proposed to foster multi-agent cooperation [8, 25].
So far, no work has examined whether causal influences on agent policies, measured via information-theoretic SVs, correspond to changes in the value function. We are the first to explore if analyzing solely the policy function provides meaningful insights into an agent’s behavior and its impact on co-players’ decisions, and if these insights align with overall team performance.
@@ 4.1 Game Model
We define intermediate state RVs $S_{t,(k)}$ and $S_{t}$ for $t$ and $k$ in the following way. Let $S_{0}=S_{0,(0)}\sim\rho$. For each $t$ we construct $S_{t,(k)}$ such that $S_{t,(k)}|S_{t,(k-1)},A_{t}^{\sigma(k)}\sim P^{\sigma(k)}$, and $S_{t+1}=S_{t,(n)}$. Each kernel $P^{i}$ has a property that it updates only state of player $i$ and the states of other agents stays fixed, i.e., $P^{i}(s_{k}|s_{k-1},a^{i})=0$ for all $j\neq i$ when $s_{k}^{j}\neq s_{k-1}^{j}$. For each $t$ a new $\sigma$ is sampled iidiid from $\mathcal{D}$.
@@
Def. 6 implies that we assume $\mathcal{G}_{s}$ to be locally Markovian, i.e., for each $i$ and step $k$, the kernel $P^{i}$ depends only on the immediately preceding state $s_{t,(k-1)}$ and the action $a_{t}^{i}$:
@@
    $P(s_{t+1}\mid s_{t},\boldsymbol{a}_{t})=\mathbb{E}_{\sigma}\sum_{\mathbf{s}\in\mathcal{S}}\prod_{k=1}^{n}P^{\sigma(k)}\left(s_{t,(k)}\mid s_{t,(k-1)},a_{t}^{\sigma(k)}\right)$ (3)
While Assumption 4 may hold only in simple environments, it enables our framework to be applied online during execution. In such cases, intermediate states $S_{t,(k)}$ can be accessed while actions are being processed. In contrast, for more complex games that support parallel execution and thereby violate Assumption 4, our SVMG framework remains applicable offline. By storing the global state trajectory $\{S_{t}\}_{t=0}^{T}$ during inference, the intermediate sequences $\{S_{t,(k)}\}_{k=1}^{n}$ can be reconstructed by Eq. (2) in a post-processing step with $\sigma$ injected externally.

### v2

@@ Abstract
To reliably deploy Multi-Agent Reinforcement Learning (MARL) systems, it is crucial to understand individual agent behaviors. While prior work typically evaluates overall team performance based on explicit reward signals, it is unclear how to infer agent contributions in the absence of any value feedback. In this work, we investigate whether meaningful insights into agent behaviors can be extracted solely by analyzing the policy distribution. Inspired by the phenomenon that intelligent agents tend to pursue convergent instrumental values, we introduce Intended Cooperation Values (ICVs), a method based on information-theoretic Shapley values for quantifying each agent’s causal influence on their co-players’ instrumental empowerment. Specifically, ICVs measure an agent’s action effect on its teammates’ policies by assessing their decision (un)certainty and preference alignment. By analyzing action effects on policies and value functions across cooperative and competitive MARL tasks, our method identifies which agent behaviors are beneficial to team success, either by fostering deterministic decisions or by preserving flexibility for future action choices, while also revealing the extent to which agents adopt similar or diverse strategies. Our proposed method offers novel insights into cooperation dynamics and enhances explainability in MARL systems.
@@ 1 Introduction
In MARL, the reward function is fundamental in shaping agent behavior, with Schmidhuber’s artificial curiosity [29] inspiring the design of intrinsic rewards to foster cooperation. As the majority of MARL algorithms (see, e.g., [26]) rely on learned value functions to estimate expected rewards in a given state, the SV naturally emerges as a credit assignment method in fully cooperative MARL scenarios. Conversely, in the absence of both value functions and reward signals, a critical question arises of how, and more importantly which, credit should be allocated, irrespective of the game type (cooperative or competitive). In this setting of missing explicit value feedback, we investigate whether SVs can be adapted to extract meaningful information about agent behavior on performance, i.e., information that remains consistent with the underlying value function, while assuming access only to the policy function.
To identify a meaningful candidate for credit, we draw on the statement by von Foerster [37]: "I shall act always so as to increase the total number of choices," thereby advocating for individual empowerment. This aligns with the instrumental convergence thesis [4], which posits that several instrumental values generally support task success and are thus pursued by many intelligent agents. Given theoretical [35] and empirical [9, 17] evidence that access to and control over a wide range of future states (choices) serves as convergently instrumental objective during training, we question whether the same holds true during inference in multi-agent settings. Concretely, is providing agents with more choices truly beneficial for task success?
@@
Figure 1: Left: a toy environment where both players must reach the target (star) to earn a shared reward, while gray cells block the path. Blue and red arrows indicate potential moves (choices). Right: probabilities (y-axis) over actions (x-axis) of P2 before (top) and after (bottom) P1 steps on the key cell and opens the green lock.
@@
Contrary to the above findings, we argue that exploiting certain situations with high certainty may be more instrumentally valuable than preserving many choices. We illustrate this concept in Fig. 1, where two agents must reach a target (star) to receive a shared reward of one. The histograms on the right depict the probability distributions over movement actions for P2 before (top) and after (bottom) P1 acted. The closed green lock cell makes P2 more likely to move downward, while still assigning some probability to the leftward action in anticipation that P1 might open the lock. Consider the case where the key opens the green lock only and P1 steps on it. This action causally affects P2’s action space by unlocking the left path. As a result, P2 not only gets more movement options, but also becomes more certain in exploiting the open path by assigning more weight toward the left action. This highlights how P2’s (increased) success is directly tied to P1’s altruistic behavior, which would correspond to an increased value under an optimal value function. In contrast, if the key were to unlock the gray locks, it would provide additional movement options that are clearly not beneficial for achieving its goal. Even if the game involves simultaneous decision-making and P2 chooses to go down, P1 would merit credit for his intention to open the path and contribute to P2’s success. In addition, we examine the impact on preference alignment between agents. To do so, we perform a state exchange intervention by placing P1 in P2’s position and v.v. to assess whether agents exhibit coordinated consensus through mutual anticipation of similar strategies. Credit is attributed to actions that enhance alignment and promote consensus.
@@
Motivated by this, we introduce an action attribution method called Intended Cooperation Value (ICV), which quantifies how cooperatively an agent behaves by measuring the causal effect of its actions on the instrumental empowerment of its teammates, using only access to their policies. By adapting causal, information-theoretic SVs, ICV uses the policy entropy to assign credit to an agent based on its intended contribution to reducing decision uncertainty and fostering preference alignment among co-players. Empirical evaluations across different settings demonstrate that ICV reliably attributes credit to behaviors that increase the likelihood of task success. Our approach provides a deeper understanding of agent interactions, particularly in scenarios where success is not explicitly measurable.
2 Related Work
There has been considerable work on understanding and explaining Reinforcement Learning (RL), primarily aimed at answering why specific actions were taken. Existing approaches include replacing the policy network with transparent models, such as decision trees [34], or incorporating auxiliary reasoning mechanisms like counterfactual analysis [41], structural causal models [23], or interpretable reward components [20]. The SV has been used as a feature attribution method in Machine Learning (ML) [11, 43], in single-agent RL [3, 32], and in MARL to assess agent contributions based on rewards [24]. Beyond explanatory purposes, SVs have also been adapted as training mechanisms in MARL, mainly to address the credit assignment problem [6, 39, 38]. In parallel, various information-theoretic intrinsic rewards have been proposed to foster multi-agent cooperation [7, 22].
So far, no work has examined whether causal influences on agent policies, measured via information-theoretic SVs, correspond to changes in the value function. We are the first to explore if analyzing solely the policy provides meaningful insights into an agent’s behavior and its impact on co-players’ decisions, and if these insights align with overall team performance to assign a reward-free credit.
3 Preliminaries
@@
Let $(\Omega,\mathcal{F},\Pr)$ be a probability space. Let $X:\Omega\rightarrow\mathcal{X}$ be a discrete Random Variable (RV) taking value $x$ in a set $\mathcal{X}$, where $p(x)=\Pr(X=x)$ is the Probability Mass Function (PMF) and $\mathcal{P}(\mathcal{X})$ is the set of probability distributions over $\mathcal{X}$. We restrict our attention to finite sets $\mathcal{X}$. The entropy of $X$ is given by $H(X)=-\sum_{x\in\mathcal{X}}p(x)\log p(x)$. For two RVs $X$ and $Y$, the mutual information is a symmetrical function and defined as $\mathrm{I}(X;Y)=H(X)-H(X\mid Y)=H(Y)-H(Y\mid X)$. For two PMFs $p$ and $q$, the Kullback–Leibler divergence is $D_{KL}(p||q)=\sum_{x\in\mathcal{X}}p(x)\log\frac{p(x)}{q(x)}$.
3.1 Multi-Agent Reinforcement Learning
@@
Interactions among $n\in\mathbb{N}$ concurrently acting agents are typically modeled using either a Markov Game (MG) [19] or an Extensive-Form Game (EFG) [25]. While in EFGs actions are selected and executed sequentially, in MGs actions are selected simultaneously but their execution may be either parallel or sequential, depending on whether the environment supports parallelization. If not, actions are executed based on a predefined or randomized agent ordering.
@@
A Markov Game is defined by the tuple
@@
where $\mathcal{N}=\{1,\dots,n\}$, with $n\in\mathbb{N}$, is the set of agents, $\mathcal{S}$ is the set of admissible states, $\mathcal{A}^{i}$ is the set of actions available to agent $i\in\mathcal{N}$, with $\boldsymbol{\mathcal{A}}=\prod_{i=1}^{n}\mathcal{A}^{i}$ being the joint action set, $P:\mathcal{S}\times\boldsymbol{\mathcal{A}}\to\mathcal{P}(\mathcal{S})$ is the state transition probability kernel, $\rho\in\mathcal{P}(\mathcal{S})$ is the initial state distribution, $R^{i}:\mathcal{S}\times\boldsymbol{\mathcal{A}}\times\mathcal{S}\to\mathbb{R}$ is the reward function of agent $i$, and $\gamma\in[0,1)$ is the discount factor controlling the effect of future rewards. At each time step $t\in\mathbb{N}$, let the RV $S_{t}$ describe the state of the MG, and the RV $A_{t}^{i}$ denote the action taken by agent $i$. Starting from an initial state $s_{0}\sim\rho$, all agents observe the current state $s_{t}\in\mathcal{S}$ and simultaneously select an action $a_{t}^{i}\in\mathcal{A}^{i}$ according to their individual policies $\pi^{i}(a\mid s_{t})$, forming a joint action $\boldsymbol{a}_{t}=(a_{t}^{1},\dots,a_{t}^{n})\in\boldsymbol{\mathcal{A}}$ drawn from the joint policy $\boldsymbol{\pi}(\boldsymbol{a}\mid s_{t})=\prod_{i=1}^{n}\pi^{i}(a\mid s_{t})$. The MG then transitions to a new state $s_{t+1}$ according to $P(s_{t+1}\mid s_{t},\boldsymbol{a}_{t})$, and each agent $i$ receives a reward $R^{i}(s_{t},\boldsymbol{a}_{t},s_{t+1})$. The performance of each agent $i$ can be measured by means of a value function defined as:
$V^{i}(s)\coloneqq\mathbb{E}_{\boldsymbol{a}_{t}\sim\boldsymbol{\pi},s_{t+1}\sim P}\left[\sum_{t=0}^{\infty}\gamma^{t}R^{i}(s_{t},\boldsymbol{a}_{t},s_{t+1})\mid s_{0}=s\right].$ (1)
Assumption 2.
We consider pre-trained Actor-Critic [15] models under the MG framework using Centralized Training with Decentralized Execution (CTDE) [21]. Further, we consider stochastic policies, which are generally more robust in real-world-like environments compared to deterministic policies [42].
@@
Assumption 1 implies that we consider domains with full state observability, though our approach can trivially be extended to partial observability and sequential decisions. Assumption 2 implies that agents act with individual policies (actors), and in fully cooperative settings, they share a global value function $V(s)$, also called centralized critic [2]. From Definition 2, we emphasize the flexibility of MGs to represent all game types [8]. We use $\boldsymbol{-i}$ to denote the set of all agent indices excluding agent $i$. We further note that the terms environment, game, and MG are used interchangeably.
@@ 3.2 Shapley Value
We build on prior work applying SVs for feature attribution in single-agent RL and ML models. Thus, let $Z$ represent the input RV.
@@
Let $\mathcal{M}=\{1,\dots,m\}$ be the set of features and $\mathcal{Z}=\prod_{i=1}^{m}\mathcal{Z}_{i}$ the input space, so that any input $z$ can be represented as an ordered set $z=\{z_{i}\mid z_{i}\in\mathcal{Z}_{i}\}_{i=1}^{m}$, with $p(z)$ being the data distribution. Then, any subset $\mathcal{K}\subseteq\mathcal{M}$ yields a partial observation $z_{\mathcal{K}}=\{z_{i}\}_{i\in\mathcal{K}}$, where $\bar{\mathcal{K}}=\mathcal{M}\setminus\mathcal{K}$ is its complement.
@@
The Shapley Value [31] is the unique method that satisfies the desirable properties of efficiency, symmetry, additivity, and null player (see Appendix A [30]). It allocates credit by computing the average marginal contribution $\Delta v(\cdot)$ of player (or feature) $i$ to the outcome of a cooperative game (or prediction) over all coalitions $\mathcal{K}\subseteq\mathcal{M}$ using a characteristic function $v:2^{\mathcal{M}}\rightarrow\mathbb{R}$, as follows:
@@ 4.1 Game Model
We define intermediate state RVs $S_{t,(k)}$ and $S_{t}$ for $t$ and $k$ in the following way. Let $S_{0}=S_{0,(0)}\sim\rho$. For each $t$, we construct $S_{t,(k)}$ such that $(S_{t,(k)}|S_{t,(k-1)},A_{t}^{\sigma(k)})\sim P^{\sigma(k)}$, and $S_{t+1}=S_{t,(n)}$. Each kernel $P^{i}$ has the property that it updates only the state of player $i$ and the states of other agents stay fixed, i.e., $P^{i}(s_{k}|s_{k-1},a^{i})=0$ for all $j\neq i$ when $s_{k}^{j}\neq s_{k-1}^{j}$. At each decision step $t$, a new order $\sigma$ is sampled iidiid from $\mathcal{D}$.
@@
Def. 6 implies that we assume $\mathcal{G}_{s}$ to be locally Markovian, i.e., for each step $k$, the next intermediate state $s_{t,(k)}$ depends only on the immediately preceding state $s_{t,(k-1)}$ and the action $a_{t}^{i}$ given that $i$ is the currently acting player:
@@
    $P(s_{t+1}\mid s_{t},\boldsymbol{a}_{t})=\mathbb{E}_{\sigma}\sum_{\mathbf{s}\in\mathcal{S}}\prod_{k=1}^{n-1}P^{\sigma(k)}\left(s_{t,(k)}\mid s_{t,(k-1)},a_{t}^{\sigma(k)}\right)$ (3)
While Assumption 4 may hold only in simple environments, it enables our framework to be applied online during execution. In such cases, intermediate states $S_{t,(k)}$ can be accessed while actions are being processed. In contrast, for more complex games that support parallel execution and thereby violate Assumption 4, our SVMG framework remains applicable offline. By storing the global state trajectory $\{S_{t}\}_{t=0}^{T}$ during inference, the intermediate sequences $\{S_{t,(k)}\}_{k=1}^{n-1}$ can be reconstructed by Eq. (2) in a post-processing step with $\sigma$ injected externally, i.e., $\sigma\sim\mathcal{D}$.